. There are five old ATM machines at Smart Bank. All probability that an ATM breaks
down in the course of a day is 0.2. If they are all in working order at the beginning a
day, and breakdowns are independent of each other, find the probability that during
the day;
i. there are no breakdowns
ii. four machines break down
iii. all machines break down
iv. at least three machines break down
1
Expert's answer
2020-04-06T15:53:48-0400

Let "X=" the number of machines which are break down: "X\sim Bin (n;p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "n=5, p=0.2"

i. there are no breakdowns 


"P(X=0)=\binom{5}{0}(0.2)^0(1-0.2)^{5-0}=""=(0.8)^5=0.32768"

ii. four machines break down 


"P(X=4)=\binom{5}{4}(0.2)^4(1-0.2)^{5-4}=""=5(0.2)^4 (0.8)=0.0064"

iii. all machines break down 


"P(X=5)=\binom{5}{5}(0.2)^5(1-0.2)^{5-5}=""=(0.2)^5=0.00032"

iv. at least three machines break down


"P(X\geq 3)=P(X=3)+P(X=4)+P(X=5)=""=\binom{5}{3}(0.2)^3(1-0.2)^{5-3}+0.0064+0.00032=""=10(0.2)^3(0.8)^2 +0.00672=0.05792"